Accelerator Update for July 26, 2006 - July 28, 2006
Wednesday July 26
The day shift began with Operations monitoring store 4857 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, and with MiniBooNE taking beam.
Operations prepared the Recycler at 9:29 AM to begin stashing antiprotons. The stash ended and Pbar resumed stacking at 9:45.
At 10:13 AM, a Linac RF station (LRF5) tripped off. It reset okay.
Pbar experts began stacking studies at 10:52 AM.
At 12:27 PM, LRF5 tripped off for the first of two times in fifty-two minutes. It reset both times.
NuMI experts began turning on their power supplies at 2:43 PM.
The Recycler began preparing to stash at 2:44 PM. The stash ended and Pbar resumed stacking at 3:36 PM.
Operations established beam to NuMI at 4:44 PM. Experts began a study period.
The Recycler began preparing to stash at 6:52 PM. The stash ended and Pbar resumed stacking at 7:13 PM.
At 7:37 PM, NuMI tripped off due to two quadrupole power supplies tripping off (QD101 & QD101F). Everything reset.
At 7:58 PM, the NuMI Lambertsons tripped off and would not reset. Operations called in a power supply expert.
The power supply expert reported at 8:16 PM that LAM61 had tripped on a Filter Rack Summation. After further investigation he reported that at least 30 capacitors were bad. At 9:15 PM, he reported that he would disconnect the filter cabinet so the experiment could run for the night. The bad capacitors will be replaced in the morning.
NuMI resumed taking beam at 9:27 PM.
At 9:44 PM, LAM61 caused the NuMI beam permit to trip, but the power supply did not trip. The permit reset.
The Recycler began preparing to stash at 11:29 PM.
At 11:59 PM, LAM60 tripped off. It reset okay.
Thursday July 27
The midnight shift began with Operations monitoring store 4857, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
At 1:45 AM, NuMI tripped off due to excessive 60hz on LAM60. It reset.
Operations terminated store 4857 at 2:02 AM. Operations began preparing CDF and D0 for access into their collision halls.
At 2:29 AM, the NuMI beam permit tripped off for the first of four times in fifteen minutes on LAM60 excessive 60hz. It reset each time.
At 3:31 AM, NuMI tripped off due to QD101. It reset okay.
CDF came out of access at 3:47 AM.
At 3:52 AM, NuMI tripped off due to excessive 60hz on a correction magnet power supply (V100). It reset.
D0 came out of access at 4:39 AM.
At 4:43 AM, NuMI tripped off due to excessive 60hz on LAM60. It reset.
Operations put the TeV into shot setup at 5:46 AM.
At 6:08 AM, NuMI tripped off for the first of two times in twenty-eight minutes due to QD101. It reset okay.
After resolving a closure problem, Operations established store 4859 at 8:11 AM with an initial luminosity of 171.2E30. This was the second best luminosity since shutdown.
The Neutron Therapy Facility (NTF) reported at 8:15 AM that they would be treating patients this afternoon.
A Controls expert reported at 9:03 AM that he had completed work on the NuMI bridge box.
The Recycler began preparing to stash at 9:07 AM. The stash ended at 9:43 AM.
At 10:59 AM, NuMI tripped off due to beam position problems. It reset okay.
At 12:29 PM, a MI power supply (South Anode) tripped off due to an interlock indication. Operations turned off the MI beam while an expert investigated. He fixed a failed resistor in MIRF3. Beam resumed at 12:57 PM.
NuMI turned off at 1:16 PM so experts could begin work on the LAM60 filters.
The Recycler began preparing to stash at 1:47 PM. The stash ended at 2:15 PM.
The LAM60 work was completed and NuMI resumed taking beam at 3:07 PM.
At 4:18 PM BRF8 tripped off on a power amplifier air flow indication. Operations contacted an RF expert. He adjusted BRF19 to compensate for BRF8 being off.
At 4:42 PM, the Booster East Anode Power Supply (EAPS) tripped off due to a problem with BRF1.
At 4:44 PM, the NuMI horn tripped off. Operations tried to contact experts, but couldn’t get a hold of anyone. Operators investigated further and discovered that the horn had slipped into a conditioning mode. Operators reset the alarm and had the horn on by 5:30 PM.
At 5:54 PM, a RF expert reported that the modulator deck power had failed for BRF1. Since replacing this modulator would hold off beam, RF techs will access the Booster to repair BRF8 during the BRF1 repair.
RF experts reported at 6:34 PM that they had replaced the air filters for BRF8.
The Booster power supply work ended. Booster came back on and Pbar, NuMI, and MiniBooNE resumed taking beam at 7:30 PM.
The Recycler began preparing to stash at 8:13 PM. The stash ended and Pbar resumed stacking at 8:43 PM.
The Recycler began preparing to stash at 11:59 PM.
Friday July 28
The midnight shift began with Operations monitoring store 4859 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
The stash ended at 12:32 AM.
At 1:02 AM, the NuMI horn tripped off on a PE1-A power supply water bit for the first of five times in about seven hours. It reset each time.
The Recycler began preparing to stash at 4:53 AM. The stash ended and Pbar resumed stacking around 5:30 AM.
The Plans for Friday and the Weekend
The plans for today are to continue monitoring store 4859. At 11 AM, the Run Coordinator will allow the TeV sector A1 pond strainer work to start as well as switching the feeders back to their normal configurations. Feeder 49 had a bad splice and was repaired yesterday. CDF and D0 will turn off during this work. The plans for the weekend are to stack, stash, and store.
Downtime: 4 hours
Path to 20 mA/hr
Current Criteria for Tevatron Shots
Machine Maintenance and Studies
Thanks for Help
Comments and Suggestions
|last modified 07/28/2006 email Fermilab|